OpenStudy (anonymous):

solve the proportion for x: 2/3=x/30

jimthompson5910 (jim_thompson5910):

2/3=x/30 2*30 = 3*x ... cross multiply 60 = 3x 3x = 60 x = ???

OpenStudy (zaarin):

X=20